# Where Giants Dwell: Probing the Environments of Early Massive Quiescent Galaxies

### v1

@@
We investigate the environments of massive quiescent galaxies at $3\lesssim z\lesssim 5$ using the GAlaxy Evolution and Assembly (GAEA) theoretical. We select galaxies with $10^{10.8}\,{\rm M}_{\sun}$ and specific star formation rate below $0.3\times t^{-1}_{\rm Hubble}$, yielding in a sample of about $5,000$ galaxies within a simulated volume of $\sim 685$ Mpc. These galaxies have formation times that cover well the range inferred from recent observational data, including a few rare objects with very short formation time-scales and early formation epochs. Model high-z quiescent galaxies are $\alpha$-enhanced and exhibit a wide range of stellar metallicity, in broad agreement with current observational estimates. Massive high-z quiescent galaxies in our model occupy a wide range of environments, from void-like regions to dense knots at the intersections of filaments. Quiescent galaxies in underdense regions typically reside in halos that collapsed early and grew rapidly at high redshift, though this trend becomes difficult to identify observationally due to large intrinsic scatter in star formation histories. The descendants of high-z massive quiescent galaxies display a broad distribution in mass and environment by $z=0$, reflecting the stochastic nature of mergers. About one-third of these systems remain permanently quenched, while most rejuvenation events are merger-driven and more common in overdense regions. Our results highlight the diversity of early quiescent galaxies and caution against assuming that all such systems trace the progenitors of present day most massive clusters.
@@ 1 Introduction
The assembly and quenching of massive galaxies are central questions in modern astrophysics. In the local Universe, where star formation can be measured accurately for large samples of galaxies using a variety of indicators, galaxies are found to follow a ‘bimodal’ distribution (e.g. Blanton et al., 2003; Kauffmann et al., 2003; Baldry et al., 2004). The latter exhibits a well defined red sequence dominated by low star formation rates, and a blue cloud populated by galaxies that are still actively forming stars at non negligible rates. The wealth of observational information available in the local Universe has allowed us to characterize how the distribution in the sSFR/color versus stellar mass plane depends on the environment, being this defined in terms of local galaxy density or through a more ‘global’ environmental proxy such as halo mass or large-scale over-density. The fraction of passive galaxies is found to be larger in over-dense regions (e.g. Kauffmann et al., 2004; Balogh et al., 2004; Baldry et al., 2006), indicating that environmental processes had a significant impact on at least a fraction of the quenched galaxy population observed today.
@@ 2 The simulation and the galaxy formation model
GAEA [^5] Details about the model, and access to a selection of data products, can be found at: https://sites.google.com/inaf.it/gaea/ is a state-of-the-art theoretical model of galaxy formation that couples prescriptions for the evolution of different baryonic components with substructure based merger trees extracted from high-resolution cosmological N-body simulations. The model builds on that presented in the original work by De Lucia & Blaizot (2007), but it has been updated significantly over the past few years. Topical developments have been:
@@
For the analysis presented below, we have selected massive quiescent galaxies at $z\gtrsim 3.1$ and $z\lesssim 5.3$. Specifically, we have considered all model galaxies in the redshift range of interest with stellar mass larger than $10^{10.8}\,{\rm M}_{\sun}$ [^6] The stellar mass cut adopted is rather arbitrary and a compromise between the large estimated stellar mass quoted in recent observational studies (e.g. Carnall et al., 2024) and the goal to have a large enough statistical sample of model galaxies. and with specific star formation rate (sSFR) lower than $0.3\times t^{-1}_{\rm Hubble}$, where $t_{\rm Hubble}$ is the age of the Universe at the corresponding redshift. Model intrinsic stellar masses have been convolved with an observational uncertainty of $0.25$ dex, which is rather conservative at these redshifts. Whenever appropriate, we will discuss in the following how results depend on this convolution. We find a total of 4788 galaxies in the entire Millennium Simulation box satisfying these criteria. We do not prevent the descendant/progenitor of the same galaxies to be counted at different snapshots. About 87 per cent of the selected galaxies are centrals and about 7 per cent are orphans, i.e. galaxies whose parent dark matter substructure has been tidally stripped below the resolution limit of the simulation. In our model, these galaxies survive for a residual merger time before merging (this is justified by the fact that baryons are much more concentrated than dark matter). Since the number density of massive quiescent galaxies decreases rapidly with increasing redshift, the sample is dominated by galaxies at the lowest end of the redshift range considered: 2195 galaxies are found at $z\sim 3.1$, 1287 at $z\sim 3.3$ and 668 at $z\sim 3.6$. There are only 12 massive quiescent galaxies at $z\sim 5.3$ - the snapshot at the highest redshift considered.
@@ 4 What is the environment of high-z quiescent galaxies?
    an ‘over-density’ estimate corresponding to the mean over-density of galaxies measured within a projected circular region of radius equal to 3 Mpc. We refer to this over-density as o33 in the following.
@@
Figure 4: Evolution of the parent halo mass, galaxy stellar mass, star formation rate, and black hole mass for the massive quiescent galaxies at $z\sim 4.5$ considered in Fig. 3. Green and magenta lines refer to the top and bottom panel respectively, with thickness of the lines and color shade increasing with over-density values as indicated in the legend included in the left panel. $\rm{M}_{200}$ and $\rm M_{BH}$} are shown for the main progenitor at each previous cosmic epoch, while for the galaxy stellar mass and SFR we show the sum of all progenitors at the corresponding epoch.
@@
Although a direct comparison is difficult, the range of environments visible in Fig. 3 is much broader from that depicted in Fig. 4 of Remus & Kimmig (2025), showing the projected surface brightness of the gas at z=3.4 with superimposed the position of the six most massive quiescent galaxies in the Magneticum Simulation Box3. In this simulation, massive quiescent galaxies at high-z do inhabit different regions of the cosmic web as is also in the case in our model, but none of them is found at the center (or close to the center) of the hottest regions (the most massive halos). As discussed in Section 1, this qualitative difference in the spatial distribution of massive quiescent galaxies is interesting because it suggests that it can be used, in principle, to constraint different physical mechanisms driving galaxy quenching at high-z.
@@ 6 What are the descendants of early quiescent galaxies?
One last question we can ask, taking advantage of our model results, is related to what are the descendants of high-z quiescent galaxies and if their future fate depends on the environments in which they reside. As noted above, about 88 per cent of the galaxies selected at high-z are centrals. Tracing their descendants down to $z=0$, we find that about 77 per cent are still centrals while about 11 per cent are orphans. Fig. 9 shows the distribution of parent halo mass (top panel) and stellar mass (middle panel) both for the galaxies selected at high redshifts (solid lines) and their descendants at $z=0$ (dashed histograms. The purple and cyan lines correspond to the galaxies that are in the extremes (15th and 85th) percentiles of the distribution of o33 at high-z. The bottom panel shows the correlation between the parent halo mass at $z=0$ and the over-density o33 at the time of observation. The figure shows clearly that the distributions of halo masses and stellar masses of the $z=0$ descendants of high-z quiescent galaxies are very broad, broader than the distribution found at the redshift of observation. This can be understood as a natural consequence of the stochasticity of merger events. The effect, that is stronger the higher the redshift, has been noted in earlier work and recognized as a generic prediction of hierarchical growth (e.g. De Lucia & Blaizot, 2007; Trenti et al., 2008; Angulo et al., 2012). Fig. 9 also shows that the large-scale overdensity has a significant impact on the fate of the high-z quiescent galaxies: galaxies that are sitting in the largest overdensities have higher probabilities to end up in a massive cluster today. There is a clear and rather strong correlation between the over-density at the time of the observation and the halo mass at $z=0$, which makes the latter a good predictor of the environment in which the descendants of high-z passive galaxies live at present. A clear but less significant difference appears when considering a local estimate of the environment. The latter is a more powerful indicator of the halo mass at the redshift of observations, but a less powerful indicator of the $z=0$ halo mass (see Fig. 14 in Appendix A). Similar results have been found in the recent work by Remus & Kimmig (2025). Their Fig. 15 shows a broad correlation between more dense environments at $z=3.4$ and larger $z=0$ halo masses. However, 75 per cent of their high-z quiescent galaxies lie in underdense environment according to their definition. As noted above, this is likely related to how AGN feedback is implemented in their simulation, and it is worth noting that their particular implementation of stellar and AGN feedback does not lead to a good agreement with the observed evolution of the galaxy stellar mass function at lower redshifts (Hirschmann et al., 2014; Steinborn et al., 2015).
@@
Figure 10: Top panels: distributions of the number of rejuvenation episodes, of the SFR during each episode, and of the associated time-scale. Bottom panels: distributions of the total number of mergers in the last snapshot just before the rejuvenation episodes, of the number of mergers with a mass ratio of at least 1 to 10, and of the associated increase of stellar mass (excluding the stellar mass that is accreted through mergers). Magenta and cyan are for the 15th and 85th percentiles of the o33 distributions.
In the framework of our model, rejuvenation is not a common event in the future lifetime of a massive quiescent galaxy at high-z. About 36 per cent of the quiescent galaxies in the selected sample remain passive down to $z=0$, experiencing no rejuvenation events. For the rest of the galaxies in the sample, the mean number of rejuvenation episodes is $\sim 1.6$ (with a median number of 1). Only very few galaxies experience a number of rejuvenation episodes larger than 4 (15 galaxies in total out of 4788). The overall fraction of galaxies experiencing rejuvenation in our model is larger than quoted in Szpila et al. (2025) but very close to those estimated in Remus & Kimmig (2025). However, we note that these numbers are very sensitive to the exact definitions adopted. The galaxies with more frequent episodes of rejuvenation events tend to reside in the less over-dense regions at the time of observations, but the number statistics are very low.
About 34 per cent of the rejuvenation episodes are associated with a star formation rate lower than $10\,{\rm M}_{\odot}\,{\rm yr}^{-1}$, but $\sim 5$ per cent of the galaxies in our sample have rejuvenation episodes with associated rates larger than $100\,{\rm M}_{\odot}\,{\rm yr}^{-1}$. A larger number of these episodes are associated with galaxies that reside in the over-dense regions at the epoch of observations. The vast majority of the rejuvenation episodes correspond to events occurring over one snapshot only of the simulation: the interval between subsequent snapshots varies between $\sim 110$ and $\sim 270$ Myr at $z\sim 5$ and $z=0$ respectively. Only a few episodes last much longer, i.e. the galaxy accretes significant amount of gas to sustain star formation over several snapshots. Rejuvenation episodes are typically associated with recent mergers. A few of the galaxies in our sample experience relatively large number of mergers, but many of these have very low mass-ratio (see bottom left and bottom middle panel). For about 15 per cent of the galaxies in the sample, the increase in mass associated with the rejuvenation episodes (subtracting the mass accreted through mergers) is non negligible (larger than $10^{10}\,{\rm M}_{\odot}$).
@@ 7 Summary and conclusions
These findings have important implications for current and future observational programs aimed at using environmental diagnostics to test quenching mechanisms or to identify the progenitors of present-day massive clusters. Both the number densities and environmental properties of massive quiescent galaxies at $z\gtrsim 3$ are expected to exhibit significant cosmic variance. Furthermore, such galaxies do not necessarily reside in the most massive halos of their epoch, nor do they always trace the sites that will evolve into the most massive clusters by $z=0$. This is only partly due to uncertainties in stellar mass estimates and largely reflects the stochastic nature of mass accretion, as discussed in previous work (e.g. De Lucia & Blaizot, 2007; Trenti et al., 2008; Overzier et al., 2009; Angulo et al., 2012, see also Fontanot et al. in preparation).
@@ Appendix A Additional figures
Figure 11: As in Fig. 3 but showing in colour only galaxies with star formation rates lower than $\sim 5\,{\rm M}_{\odot}\,{\rm yr}^{-1}$.
Figure 12: As in Fig. 3 but showing in colour only galaxies with star formation rates larger than $\sim 80\,{\rm M}_{\odot}\,{\rm yr}^{-1}$.
@@
Fig. 13 is the equivalent of Fig. 7 but showing the average star formation histories of galaxies in our quiescent sample in the 25th (dark cyan) and 75th (dark magenta) percentiles of the distribution of o33. While there is a clear different (albeit with a very large scatter) between the extremes considering the n33 distribution as discussed in the main body of the paper, this difference is largely removed when considering larger scale over-densities.

### v2

@@
We investigate the environments of massive quiescent galaxies at $3\lesssim z\lesssim 5$ using the GAlaxy Evolution and Assembly (GAEA) theoretical model. We select galaxies with $10^{10.8}\,{\rm M}_{\sun}$ and specific star formation rate below $0.3\times t^{-1}_{\rm Hubble}$, yielding in a sample of about $5,000$ galaxies within a simulated volume of $\sim 685$ Mpc. These galaxies have formation times that cover well the range inferred from recent observational data, including a few rare objects with very short formation time-scales and early formation epochs. Model high-z quiescent galaxies are $\alpha$-enhanced and exhibit a wide range of stellar metallicity, in broad agreement with current observational estimates. Massive high-z quiescent galaxies in our model occupy a wide range of environments, from void-like regions to dense knots at the intersections of filaments. Quiescent galaxies in under-dense regions typically reside in halos that collapsed early and grew rapidly at high redshift, though this trend becomes difficult to identify observationally due to large intrinsic scatter in star formation histories. The descendants of high-z massive quiescent galaxies display a broad distribution in mass and environment by $z=0$, reflecting the stochastic nature of mergers. About one-third of these systems remain permanently quenched, while most rejuvenation events are merger-driven and more common in overdense regions. Our results highlight the diversity of early quiescent galaxies and caution against assuming that all such systems trace the progenitors of present day most massive clusters.
@@ 1 Introduction
The assembly and quenching of massive galaxies are central questions in modern astrophysics. In the local Universe, where star formation can be measured accurately for large samples of galaxies using a variety of indicators, galaxies are found to follow a ‘bimodal’ distribution (e.g. Blanton et al., 2003; Kauffmann et al., 2003; Baldry et al., 2004). This bimodality exhibits a well defined red sequence dominated by low star formation rates, and a blue cloud populated by galaxies that are still actively forming stars at non negligible rates. The wealth of observational information available in the local Universe has allowed us to characterize how the distribution in the sSFR/color versus stellar mass plane depends on the environment, defined either in terms of local galaxy density or through a more ‘global’ environmental proxy such as halo mass or large-scale over-density. The fraction of passive galaxies is found to be larger in over-dense regions (e.g. Kauffmann et al., 2004; Balogh et al., 2004; Baldry et al., 2006), indicating that environmental processes had a significant impact on at least a fraction of the quenched galaxy population observed today.
@@ 2 The simulation and the galaxy formation model
GAEA [^5] Details about the model, and access to a selection of data products, can be found at: https://sites.google.com/inaf.it/gaea/ is a state-of-the-art theoretical model of galaxy formation that couples prescriptions for the evolution of different baryonic components with substructure based merger trees extracted from high-resolution cosmological N-body simulations. The model builds on that presented in the original work by De Lucia and Blaizot (2007), but it has been updated significantly over the past few years. Topical developments have been:
@@
For the analysis presented below, we have selected massive quiescent galaxies at $z\gtrsim 3.1$ and $z\lesssim 5.3$. Specifically, we have considered all model galaxies in the redshift range of interest with stellar mass larger than $10^{10.8}\,{\rm M}_{\sun}$ [^6] The stellar mass cut adopted is rather arbitrary and a compromise between the large estimated stellar mass quoted in recent observational studies (e.g. Carnall et al., 2024) and the goal to have a large enough statistical sample of model galaxies., after accounting for an observational uncertainty of 0.25 dex, and with specific star formation rate (sSFR) lower than $0.3\times t^{-1}_{\rm Hubble}$, where $t_{\rm Hubble}$ is the age of the Universe at the corresponding redshift. We note that the assumed uncertainty is rather conservative at these redshift, and that a different selection would not change qualitatively the results presented in this work. Whenever appropriate, we will discuss in the following how results depend on this convolution. We find a total of 4788 galaxies in the entire Millennium Simulation box satisfying these criteria. We do not prevent the descendant/progenitor of the same galaxies to be counted at different snapshots. About 87 per cent of the selected galaxies are centrals and about 7 per cent are orphans, i.e. galaxies whose parent dark matter substructure has been tidally stripped below the resolution limit of the simulation. In our model, these galaxies survive for a residual merger time before merging (this is justified by the fact that baryons are much more concentrated than dark matter). Since the number density of massive quiescent galaxies decreases rapidly with increasing redshift, the sample is dominated by galaxies at the lowest end of the redshift range considered: 2195 galaxies are found at $z\sim 3.1$, 1287 at $z\sim 3.3$ and 668 at $z\sim 3.6$. There are only 12 massive quiescent galaxies at $z\sim 5.3$ - the snapshot at the highest redshift considered.
@@ 4 What is the environment of high-z quiescent galaxies?
    an ‘over-density’ estimate corresponding to the mean over-density of galaxies measured within a projected circular region of radius equal to 3 Mpc. This is normalized to the average value obtained at each redshift, by randomly drawing 1000 spheres with radius equal to 3 Mpc in the entire volume of the simulation. We refer to this over-density as o33 in the following.
@@
Figure 4: Evolution of the parent halo mass, galaxy stellar mass, star formation rate, and black hole mass for the massive quiescent galaxies at $z\sim 4.5$ considered in Fig. 3. Green and magenta lines refer to the top and bottom panel respectively, with thickness of the lines and color shade increasing with over-density values as indicated in the legend included in the left panel. Different line styles are used to show different galaxies for clearer visualization. $\rm{M}_{200}$ and $\rm M_{BH}$} are shown for the main progenitor at each previous cosmic epoch, while for the galaxy stellar mass and SFR we show the sum of all progenitors at the corresponding epoch.
@@
Although a direct comparison is difficult, the range of environments visible in Fig. 3 is much broader from that depicted in Fig. 4 of Remus and Kimmig (2025), showing the projected surface brightness of the gas at z=3.4 with superimposed the position of the six most massive quiescent galaxies in the Magneticum Simulation Box3. In this simulation, massive quiescent galaxies at high-z do inhabit different regions of the cosmic web as is also in the case in our model, but none of them is found at the center (or close to the center) of the hottest regions (the most massive halos). As discussed in Section 1, this qualitative difference in the spatial distribution of massive quiescent galaxies is interesting because it suggests that it can be used, in principle, to constraint different physical mechanisms driving galaxy quenching at high-z.
@@ 6 What are the descendants of early quiescent galaxies?
One last question we can ask, taking advantage of our model results, is related to what are the descendants of high-z quiescent galaxies and if their future fate depends on the environments in which they reside. As noted above, about 88 per cent of the galaxies selected at high-z are centrals. Tracing their descendants down to $z=0$, we find that about 77 per cent are still centrals while about 11 per cent are orphans. Fig. 9 shows the distribution of parent halo mass (top panel) and stellar mass (middle panel) both for the galaxies selected at high redshifts (solid lines) and their descendants at $z=0$ (dashed histograms). The purple and cyan lines correspond to the galaxies that are in the extremes (15th and 85th) percentiles of the distribution of o33 at high-z. The bottom panel shows the correlation between the parent halo mass at $z=0$ and the over-density o33 at the time of observation. The figure shows clearly that the distributions of halo masses and stellar masses of the $z=0$ descendants of high-z quiescent galaxies are very broad, broader than the distribution found at the redshift of observation. This can be understood as a natural consequence of the stochasticity of merger events. The effect, that is stronger the higher the redshift, has been noted in earlier work and recognized as a generic prediction of hierarchical growth (e.g. De Lucia and Blaizot, 2007; Trenti et al., 2008; Angulo et al., 2012). Fig. 9 also shows that the large-scale overdensity has a significant impact on the fate of the high-z quiescent galaxies: galaxies that are sitting in the largest overdensities have higher probabilities to end up in a massive cluster today. There is a clear and rather strong correlation between the over-density at the time of the observation and the halo mass at $z=0$, which makes the latter a good predictor of the environment in which the descendants of high-z passive galaxies live at present. A clear but less significant difference appears when considering a local estimate of the environment. The latter is a more powerful indicator of the halo mass at the redshift of observations, but a less powerful indicator of the $z=0$ halo mass (see Fig. 14 in Appendix A). Similar results have been found in the recent work by Remus and Kimmig (2025). Their Fig. 15 shows a broad correlation between more dense environments at $z=3.4$ and larger $z=0$ halo masses. However, 75 per cent of their high-z quiescent galaxies lie in underdense environment according to their definition. As noted above, this is likely related to how AGN feedback is implemented in their simulation, and it is worth noting that their particular implementation of stellar and AGN feedback does not lead to a good agreement with the observed evolution of the galaxy stellar mass function at lower redshifts (Hirschmann et al., 2014; Steinborn et al., 2015).
@@
Figure 10: Top panels: distributions of the number of rejuvenation episodes, of the SFR during each episode, and of the associated time-scale. Bottom panels: distributions of the total number of mergers in the last snapshot just before the rejuvenation episodes, of the number of mergers with a mass ratio of at least 1 to 10, and of the associated fractional increase of stellar mass (excluding the stellar mass that is accreted through mergers). Magenta and cyan are for the 15th and 85th percentiles of the o33 distributions.
In the framework of our model, rejuvenation is not a common event in the future lifetime of a massive quiescent galaxy at high-z. About 36 per cent of the quiescent galaxies in the selected sample remain passive down to $z=0$, experiencing no rejuvenation events. For the rest of the galaxies in the sample, the mean number of rejuvenation episodes is $\sim 1.6$ (with a median number of 1). Only very few galaxies experience a number of rejuvenation episodes larger than 4 (15 galaxies in total out of 4788). The overall fraction of galaxies experiencing rejuvenation in our model is larger than quoted in Szpila et al. (2025) but very close to those estimated in Remus and Kimmig (2025). However, we note that these numbers are very sensitive to the exact definitions adopted. The galaxies with more frequent episodes of rejuvenation events tend to reside in the less over-dense regions at the time of observations, but the number statistics are very low.
About 34 per cent of the rejuvenation episodes are associated with a star formation rate lower than $10\,{\rm M}_{\odot}\,{\rm yr}^{-1}$, but $\sim 5$ per cent of the galaxies in our sample have rejuvenation episodes with associated rates larger than $100\,{\rm M}_{\odot}\,{\rm yr}^{-1}$. A larger number of these episodes are associated with galaxies that reside in the over-dense regions at the epoch of observations. The vast majority of the rejuvenation episodes correspond to events occurring over one snapshot only of the simulation: the interval between subsequent snapshots varies between $\sim 110$ and $\sim 270$ Myr at $z\sim 5$ and $z=0$ respectively. Only a few episodes last much longer, i.e. the galaxy accretes significant amount of gas to sustain star formation over several snapshots. Rejuvenation episodes are typically associated with recent mergers. A few of the galaxies in our sample experience relatively large number of mergers, but many of these have very low mass-ratio (see bottom left and bottom middle panel). For about 15 per cent of the galaxies in the sample, the increase in mass associated with the rejuvenation episodes (subtracting the mass accreted through mergers) is non negligible (larger than $10^{10}\,{\rm M}_{\odot}$). The fractional mass increase with respect to the present stellar mass tends to be larger for galaxies in underdense environments (bottom-right panel), due to the large difference between the present day stellar mass of galaxies residing in overdense and underdense environments (see middle panel of Fig. 9).
@@ 7 Summary and conclusions
These findings have important implications for current and future observational programs aimed at using environmental diagnostics to test quenching mechanisms or to identify the progenitors of present-day massive clusters. Both the number densities and environmental properties of massive quiescent galaxies at $z\gtrsim 3$ are expected to exhibit significant cosmic variance. Furthermore, such galaxies do not necessarily reside in the most massive halos of their epoch, nor do they always trace the sites that will evolve into the most massive clusters by $z=0$. This is only partly due to uncertainties in stellar mass estimates and largely reflects the stochastic nature of mass accretion, as discussed in previous work (e.g. De Lucia and Blaizot, 2007; Trenti et al., 2008; Overzier et al., 2009; Angulo et al., 2012, see also Fontanot et al. in preparation).
@@ Appendix A Additional figures
Figure 11: As in Fig. 3 but showing in color only galaxies with star formation rates lower than $\sim 5\,{\rm M}_{\odot}\,{\rm yr}^{-1}$.
Figure 12: As in Fig. 3 but showing in color only galaxies with star formation rates larger than $\sim 80\,{\rm M}_{\odot}\,{\rm yr}^{-1}$.
@@
Fig. 13 is the equivalent of Fig. 7 but showing the average star formation histories of galaxies in our quiescent sample in the 25th (dark cyan) and 75th (dark magenta) percentiles of the distribution of o33. While there is a clear different (albeit with a very large scatter) between the extremes considering the n33 distribution as discussed in the main body of the paper, this difference is largely removed when considering larger scale overdensities.